A pharmacist wants to mix a 6% solution with 5 liters of a 12% solution to get a mixture that is an 8% solution. How much of the 6% solution should he use?
Answer by TimothyLamb(4379)

---
12*5 + 6v = 8(v + 5)
60 + 6v = 8v + 40
2v = 20
---
v = 10 liters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php